OpenStudy (anonymous):

how do you solve this?? o_0? the cube root of 5x+7 +3=1

OpenStudy (anonymous):



OpenStudy (anonymous):

subtract 3 from both sides and then cube it

OpenStudy (anonymous):



OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

what do u do to the -3+1??

OpenStudy (anonymous):

that would be -2 (-2)^3=-8

OpenStudy (anonymous):

i mean to the -2?

OpenStudy (anonymous):

cube it

OpenStudy (anonymous):

thanks =D

OpenStudy (anonymous):

:) caN u show me the exact steps?? i got -3 in the book

OpenStudy (anonymous):

yup =D

OpenStudy (turingtest):

1)subtract 3 from both sides now what do you have?

OpenStudy (turingtest):

@soccerdoggie6 please do not give the complete answer, the goal is to get them to figure it out themselves

OpenStudy (turingtest):

otherwise they just copy the work...

OpenStudy (turingtest):

so we need to make sure they are trying :)

OpenStudy (turingtest):

so continue to engage the student and you will earn more points @soccerdoggie6

OpenStudy (anonymous):

well turing test did you see the work above that thats what i was doing and the points dont matter

OpenStudy (anonymous):

yea thanks i dont see any more im stuck :(

OpenStudy (turingtest):

well, by the code of conduct you are not supposed to just print the entire answer, you are suposed to teach, so if you do that I will have to dock you points :(

OpenStudy (turingtest):

@cos00155079 try each step individually 1)subtract 3 from each side please write what you have after you do that

OpenStudy (anonymous):

i showed the work if the person looking at the work doesnt want to learn thats their own problem. and i didnt give the answer right away. go ahead and subtract points what do i care

OpenStudy (anonymous):

ok ok stop fighting just help me please i have 5x+7=-2??? do u subtract the 7?

OpenStudy (turingtest):

no, you cube it as soccerdoggie said do you know how to cube -2 ? i.e. what is (-2)^3 ?

OpenStudy (anonymous):

8

OpenStudy (turingtest):

close, but actually it's negative (-2)(-2)(-2)=4(-2)=-8 remember negative times negative is positive, and positive times negative is negative

OpenStudy (turingtest):

so what is the expression now?

OpenStudy (turingtest):

...the whole equation I mean

OpenStudy (anonymous):

sorry my internet failed...i have 5x+7=-8

OpenStudy (turingtest):

excellent, now subtract 7 from each side what do you have next?

OpenStudy (anonymous):

i solved it... 5x=-15

OpenStudy (anonymous):

then i divided by 5 and got the answer =-3

OpenStudy (turingtest):

very nice :)